# Contact graphs of boxes with unidirectional contactsthanks: This research is partially supported by the ANR GATO, under contract ANR-16-CE40-0009 and ANR project GRALMECO (ANR-21-CE48-0004-01)

### v3

@@ 7 Recognition
Computing the boxicity of a bipartite graph is a difficult problem. It is known that deciding whether a bipartite graph has boxicity two is NP-complete [24]. Furthermore, it is proven in [1] that it is not possible to approximate the boxicity of a bipartite graph within a $O(n^{0.5-\varepsilon})$-factor in polynomial time, unless $NP=ZPP$. By Lemma 20, for every bipartite graph $B$ there is a graph $B^{\prime}$ (obtained in polynomial time) such that the minimum value $d$ for which $B^{\prime}$ belongs to $d$-CBU is exactly $d=box(B)+1$.
Corollary 28
@@
By Lemma 20, this problem would be solved, for $d\geq 3$, if the following problem admits a positive answer.
Problem 30
@@ 7.1 Recognition through forbidden induced subgraphs
As CBU and $d$-CBU are closed under induced subgraphs, they are characterized by a set of minimal excluded induced subgraphs, $\mathcal{F}_{CBU}$ and $\mathcal{F}_{d-CBU}$. If one of these sets is finite, then recognizing the corresponding class becomes polynomial-time tractable. Thus by Corollary 28, the set $\mathcal{F}_{3-CBU}$ (resp. $\mathcal{F}_{d-CBU}$ for $d\geq 4$) is not finite, unless $P=NP$ (resp. unless $NP=ZPP$). For the set $\mathcal{F}_{2-CBU}$ (resp. $\mathcal{F}_{CBU}$), we are sure that it is infinite. Indeed, Theorem 14 (resp. Theorem 26) provides an infinite sequence of graphs $\left(G_{i}\right)_{i\geq 0}$ not in 2-CBU (resp. not in CBU) such that the girth of $G_{i}$ is at least $i$. If there was an $n$ such that every graph in $\mathcal{F}_{2-CBU}$ (resp. $\mathcal{F}_{CBU}$) has at most $n$ vertices, then to exclude $G_{n+1}$ one would need to have a tree in $\mathcal{F}_{2-CBU}$ (resp. $\mathcal{F}_{CBU}$). This is not the case as for every tree $T$, we have that $T\in 2$-$CBU\subseteq CBU$.
@@ 7.2 Recognition through homogeneous arc labelings
By Theorem 24, a graph $G$ belongs to CBU if and only if it admits a homogeneous arc labeling. If we are given an orientation of a graph $G$ it is simple to check whether this orientation admits such labeling. For example, one can use linear programming. For each arc $uv$, set a variable $\ell_{uv}$ corresponding to a label, and for any two incident arcs, add a constraint. For two arcs $uv$ and $uw$ (resp. $uv$ and $wv$), the constraint is $\ell_{uv}=\ell_{uw}$ (resp. $\ell_{uv}=\ell_{wv}$). For two arcs $uv$ and $vw$, the constraint is $\ell_{uv}+1\leq\ell_{vw}$. Problem 31 thus reduces to deciding whether a graph $G$ admits an orientation that is homogeneously labelable. In the following we characterize such orientations.
@@
Proof. For the "only if" part, consider a badly oriented cycle $(v_{0},v_{1},\ldots,v_{n-1})$ with arcs $v_{n-1}v_{0}$ and $v_{0}v_{1}$, but with no vertex $v_{j}\neq v_{0}$ whose incident arcs are $v_{j+1}v_{j}$ and $v_{j}v_{j-1}$. This latter condition implies that in any homogeneous labeling the sequence of labels for the edges (without considering their orientation) $v_{0}v_{1},v_{1}v_{2},\ldots,v_{n-2}v_{n-1},v_{n-1}v_{0}$ is non-decreasing, while the former condition implies that the label of $v_{0}v_{1}$ is greater than the one of $v_{n-1}v_{0}$, a contradiction. Thus this orientation of $G$ does not allow any homogeneous labeling.
@@
Proof. If $G^{\prime}$ had a badly oriented cycle $C$, this one should go through a newly added arc $u^{\prime}v_{j}$. If $v_{i}\notin C$, by replacing the arc $u^{\prime}v_{j}$ by the path $(u^{\prime},v_{i},u,v_{j})$ one would obtain a badly oriented cycle in $G$, a contradiction. We thus assume that $v_{i}\notin C$, and now by replacing the arc $u^{\prime}v_{j}$ by the path $(u^{\prime},v_{i},u,v_{j})$ we obtain a badly oriented closed walk $W$ (that is a walk where there are consecutive "forward" arcs, but no consecutive "backward" arcs). Let us denote $P$ and $P^{\prime}$ the sub-paths of $C\setminus\{u^{\prime}v_{j}\}\subsetneq G$ linking $v_{i}$ and $v_{j}$, and linking $u^{\prime}$ and $v_{i}$, respectively.
@@
This concludes the proof of the claim $\Box$
So now, by recurrence on the number of vertices we can assume that $G^{\prime}$ has a homogeneous labeling, and let $\ell$ be the label of the arcs outgoing from $u^{\prime}$. In that case one can derive a labeling of $G$ by keeping the same labels, and by setting the label $\ell$ for the arcs outgoing from $u$. It is easy to check that this labeling is homogeneous. $\Box$
Note that Theorem 33 provides another proof that CBU contains every bipartite graph. Indeed, orienting all the edges from one part toward the other, the direction of the arcs alternate along any cycle, and so there is no badly oriented cycle. Actually, we can go a little further.
Theorem 35
@@
Proof. A graph $G$ with circular chromatic number $\chi_{c}(G)\leq 5/2$ has a homomorphism into the circular complete graph $K_{5/2}$ that is the 5-cycle. As this graph belongs to CBU the theorem follows from Theorem 37. $\Box$
Note that we cannot replace $5/2$ by $8/3$ in Theorem 35, as one can easily check that every orientation of $K_{8/3}$ contains a badly oriented cycle.
Problem 36
@@
Proof. By Theorem 24, the graph $H$ admits a homogeneous arc labeling, $\ell_{H}$. Orient the edges of $G$ in such a way that $uv\in E(G)$ is oriented as the edge $\gamma(u)\gamma(v)\in E(H)$, that is from $u$ to $v$ if and only if $\gamma(u)\gamma(v)$ is oriented from $\gamma(u)$ to $\gamma(v)$ in $H$. Similarly we copy the labeling of $H$’s arcs by setting $\ell_{G}(uv)=\ell_{H}(\gamma(u)\gamma(v))$. One can easily check that this is a homogeneous arc labeling of $G$, and thus that $G$ belongs to CBU. $\Box$
@@ 8 Chromatic Number and Independent Sets
Theorem 38 (Magnant and Martin [26])
@@
Proof. The first statement follows from Grötzsch’s theorem. The second statement follows from graphs constructed by Jones [22], which were proved to have independence number $\alpha(G)=(n+1)/3$. Those graphs form a sequence $J_{1},J_{2},\ldots$ such that $J_{1}$ is the 5-cycle $(a_{1},b_{1},c_{1},d,e)$, and such that $J_{i+1}$ is obtained from $J_{i}$ by adding three vertices $a_{i+1},b_{i+1},c_{i+1}$ such that $N(a_{i+1})=\{b_{i},b_{i+1}\}$, $N(b_{i+1})=\{a_{i+1},c_{i+1}\}$, and $N(c_{i+1})=\{a_{i},c_{i},b_{i+1}\}$ (see Figure 8). It is already known that those graphs are planar, and it does only remain to show that they belong to CBU. Let us do so by exhibiting a homogeneous arc labeling $\ell$. This labeling is such that for any $i\geq 1$ we orient the edges $a_{i}b_{i}$ and $b_{i}c_{i}$ toward $b_{i}$, we orient the edges $x_{i}y_{i+1}$, for $x,y\in\{a,b,c\}$, from $x_{i}$ towards $y_{i+1}$, and we set $\ell(a_{i}b_{i})=\ell(a_{i}c_{i+1)}=2i$, $\ell(c_{i}b_{i})=\ell(c_{i}c_{i+1)}=2i$, and $\ell(b_{i}a_{i+1)}=2i+1$. By examining Figure 8 it is clear that this is a homogeneous arc labeling. $\Box$
@@
A positive answer to this question, would give support to two conjectures. Let $\mathcal{P}_{g\geq 5}$ be the set of planar graph with girth at least five, and let $\mathcal{P}^{f}_{g\geq 4}$ be the set of planar graph with girth at least four, where every 4-cycle bounds a face. Clearly $\mathcal{P}_{g\geq 5}\subsetneq\mathcal{P}^{f}_{g\geq 4}$, since these classes avoid Jones graphs it is conjectured that graphs in $\mathcal{P}_{g\geq 5}$, or more generally graphs in $\mathcal{P}^{f}_{g\geq 4}$, have fractional chromatic number at most $c$, for some $c<3$ [13, 15]. However, our problem is not a sub-case of these conjectures (as $K_{2,t}$ belongs to $2$-CBU $\setminus\ \mathcal{P}^{f}_{g\geq 4}$), nor a super-case (as $\mathcal{P}_{g\geq 5}\setminus 2$-CBU is not empty, by Theorem 14).
@@ 9 Computational hardness for many problems
We have seen (c.f. Theorem 8, Corollary 9, and Corollary 10) that many 1-subdivided graphs belong to CBU, or even to 3- or 4-CBU. For $(\geq 2)$-subdivided graphs, the picture is even simpler.
Theorem 42
@@
Proof. Let us denote $v_{1},\ldots,v_{n}$ the vertices of $G$, and let $m=|E(G)|$. To construct a CBU representation for any $(\geq 2)$-subdivision, we start by assigning each vertex $v_{i}$ to the box $[3i,3i+1]\times[n-i,n-i+1]\times[0,2m]$. Then consider each edge $e$ of $G$ in any given order. For the $k^{\text{th}}$ edge $e$ assume it links $v_{i}$ and $v_{j}$, for some $i<j$, and assume $e$ is replaced by the path $(v_{i},u_{1},\ldots,u_{r},v_{j})$ for some $r\geq 2$. Here, $u_{1}$ is assigned to $[3i+1,3i+2]\times[n-j,n-i+1]\times[2k-1,2k]$, while the vertices $u_{\ell}$ with $2\leq\ell\leq r$ are assigned to $[3i+2+(\ell-2)(3j-3i-2)/(r-1),3i+2+(\ell-1)(3j-3i-2)/(r-1)]\times[n-j,n-j+1]\times[2k-1,2k]$ (see Figure 9). One can easily check that the obtained representation is a 3-CBU representation of the subdivided graph. $\Box$
@@
Proof. For Minimum Feedback Vertex Set and Cutwidth, this follows from the fact that these problems are NP-hard, and that for any instance, subdividing an edge does not change the solution. For Maximum Cut, it follows from its APX-hardness and the fact that the maximum cut of a graph $G$ and its 2-subdivision $G_{2\text{-sub}}$ verify $mc(G)=mc(G_{2\text{-sub}})-2|E(G)|$ and $3|E(G)|/2=|E(G_{2\text{-sub}})|/2\leq mc(G_{2\text{-sub}})\leq|E(G_{2\text{-sub}})|=3|E(G)|$. The other problems are shown APX-hard even when restricted to $6$-subdivided graphs [6]. $\Box$
@@
Proof. As these problems belong to NP, it remains to show that they are NP-hard for 2-CBU graphs. Let us first show that the induced subgraphs of grids (so called grid graphs) belong to 2-CBU. Consider the $n\times n$ grid $G$ such that $V(G)=\{1,\ldots,n\}\times\{1,\ldots,n\}$, and such that the neighbors of any vertex $(i,j)$ are $\{(i,j-1)(i-1,j),(i,j+1),(i+1,j)\}\cap\{1,\ldots,n\}\times\{1,\ldots,n\}$. Since it suffices to delete some boxes to obtain an induced subgraph, the claim follows by constructing a 2-CBU representation for any such grid $G$. This construction is obtained by mapping any vertex $(i,j)$ to the box $[i+j-1,i+j]\times[2i-2j,2i-2j+3]$ (see Figure 10). As Domination [7], Hamiltonian Path, and Hamiltonian cycle [21] are NP-hard for grid graphs, those problems are NP-hard for 2-CBU graphs.
@@
For the problems Maximum Independent Set and Minimum Vertex Cover, we have to consider a variant of grid graphs, the graph $R^{\prime}(n_{1},n_{2})$ depicted in Figure 11, and it is easy to see how to modify the construction above in order to obtain a 2-CBU representation of this type of graphs. Again, this implies that every induced subgraph of such a graph belongs to 2-CBU. As the problems Maximum Independent Set and Minimum Vertex Cover are NP-hard for this class (see the proof of Theorem 10 in [19]), those problems are NP-hard for 2-CBU graphs $\Box$
Refer to caption$(i-1,j)$$(i,j-1)$$(i+1,j)$$(i,j+1)$$(i-1,j+1)$$(i+1,j-1)$

### v4

@@ 7 Recognition
Computing the boxicity of a bipartite graph is a difficult problem. It is known that deciding whether a bipartite graph has boxicity two is NP-complete [24]. Furthermore, it is proven in [1] that it is not possible to approximate the boxicity of a bipartite graph within a $O(n^{0.5-\varepsilon})$-factor in polynomial time, unless $NP=ZPP$. By Lemma 6.26, for every bipartite graph $B$ there is a graph $B^{\prime}$ (obtained in polynomial time) such that the minimum value $d$ for which $B^{\prime}$ belongs to $d$-CBU is exactly $d=box(B)+1$.
Corollary 7.40.
@@
By Lemma 6.26, this problem would be solved, for $d\geq 3$, if the following problem admits a positive answer.
Problem 7.42.
@@ 7.1 Recognition through forbidden induced subgraphs
As CBU and $d$-CBU are closed under induced subgraphs, they are characterized by a set of minimal excluded induced subgraphs, $\mathcal{F}_{CBU}$ and $\mathcal{F}_{d-CBU}$. If one of these sets is finite, then recognizing the corresponding class becomes polynomial-time tractable. Thus by Corollary 7.40, the set $\mathcal{F}_{3-CBU}$ (resp. $\mathcal{F}_{d-CBU}$ for $d\geq 4$) is not finite, unless $P=NP$ (resp. unless $NP=ZPP$). For the set $\mathcal{F}_{2-CBU}$ (resp. $\mathcal{F}_{CBU}$), we are sure that it is infinite. Indeed, Theorem 5.18 (resp. Theorem 6.37) provides an infinite sequence of graphs $\left(G_{i}\right)_{i\geq 0}$ not in 2-CBU (resp. not in CBU) such that the girth of $G_{i}$ is at least $i$. If there was an $n$ such that every graph in $\mathcal{F}_{2-CBU}$ (resp. $\mathcal{F}_{CBU}$) has at most $n$ vertices, then to exclude $G_{n+1}$ one would need to have a tree in $\mathcal{F}_{2-CBU}$ (resp. $\mathcal{F}_{CBU}$). This is not the case as for every tree $T$, we have that $T\in 2$-$CBU\subseteq CBU$.
@@ 7.2 Recognition through homogeneous arc labelings
By Theorem 6.34, a graph $G$ belongs to CBU if and only if it admits a homogeneous arc labeling. If we are given an orientation of a graph $G$ it is simple to check whether this orientation admits such labeling. For example, one can use linear programming. For each arc $uv$, set a variable $\ell_{uv}$ corresponding to a label, and for any two incident arcs, add a constraint. For two arcs $uv$ and $uw$ (resp. $uv$ and $wv$), the constraint is $\ell_{uv}=\ell_{uw}$ (resp. $\ell_{uv}=\ell_{wv}$). For two arcs $uv$ and $vw$, the constraint is $\ell_{uv}+1\leq\ell_{vw}$. Problem 7.43 thus reduces to deciding whether a graph $G$ admits an orientation that is homogeneously labelable. In the following we characterize such orientations.
@@
Proof 7.46.
For the "only if" part, consider a badly oriented cycle $(v_{0},v_{1},\ldots,v_{n-1})$ with arcs $v_{n-1}v_{0}$ and $v_{0}v_{1}$, but with no vertex $v_{j}\neq v_{0}$ whose incident arcs are $v_{j+1}v_{j}$ and $v_{j}v_{j-1}$. This latter condition implies that in any homogeneous labeling the sequence of labels for the edges (without considering their orientation) $v_{0}v_{1},v_{1}v_{2},\ldots,v_{n-2}v_{n-1},v_{n-1}v_{0}$ is non-decreasing, while the former condition implies that the label of $v_{0}v_{1}$ is greater than the one of $v_{n-1}v_{0}$, a contradiction. Thus this orientation of $G$ does not allow any homogeneous labeling.
@@
Proof 7.47.
If $G^{\prime}$ had a badly oriented cycle $C$, this one should go through a newly added arc $u^{\prime}v_{j}$. If $v_{i}\notin C$, by replacing the arc $u^{\prime}v_{j}$ by the path $(u^{\prime},v_{i},u,v_{j})$ one would obtain a badly oriented cycle in $G$, a contradiction. We thus assume that $v_{i}\notin C$, and now by replacing the arc $u^{\prime}v_{j}$ by the path $(u^{\prime},v_{i},u,v_{j})$ we obtain a badly oriented closed walk $W$ (that is a walk where there are consecutive "forward" arcs, but no consecutive "backward" arcs). Let us denote $P$ and $P^{\prime}$ the sub-paths of $C\setminus\{u^{\prime}v_{j}\}\subsetneq G$ linking $v_{i}$ and $v_{j}$, and linking $u^{\prime}$ and $v_{i}$, respectively.
@@
This concludes the proof of the claim
So now, by recurrence on the number of vertices we can assume that $G^{\prime}$ has a homogeneous labeling, and let $\ell$ be the label of the arcs outgoing from $u^{\prime}$. In that case one can derive a labeling of $G$ by keeping the same labels, and by setting the label $\ell$ for the arcs outgoing from $u$. It is easy to check that this labeling is homogeneous.
Note that Theorem 7.45 provides another proof that CBU contains every bipartite graph. Indeed, orienting all the edges from one part toward the other, the direction of the arcs alternate along any cycle, and so there is no badly oriented cycle. Actually, we can go a little further.
Theorem 7.48.
@@
Proof 7.49.
A graph $G$ with circular chromatic number $\chi_{c}(G)\leq 5/2$ has a homomorphism into the circular complete graph $K_{5/2}$ that is the 5-cycle. As this graph belongs to CBU the theorem follows from Theorem 7.51.
Note that we cannot replace $5/2$ by $8/3$ in Theorem 7.48, as one can easily check that every orientation of $K_{8/3}$ contains a badly oriented cycle.
Problem 7.50.
@@
Proof 7.52.
By Theorem 6.34, the graph $H$ admits a homogeneous arc labeling, $\ell_{H}$. Orient the edges of $G$ in such a way that $uv\in E(G)$ is oriented as the edge $\gamma(u)\gamma(v)\in E(H)$, that is from $u$ to $v$ if and only if $\gamma(u)\gamma(v)$ is oriented from $\gamma(u)$ to $\gamma(v)$ in $H$. Similarly we copy the labeling of $H$’s arcs by setting $\ell_{G}(uv)=\ell_{H}(\gamma(u)\gamma(v))$. One can easily check that this is a homogeneous arc labeling of $G$, and thus that $G$ belongs to CBU.
@@ 8 Chromatic Number and Independent Sets
Theorem 8.53 (Magnant and Martin [26]).
@@
Proof 8.56.
The first statement follows from Grötzsch’s theorem. The second statement follows from graphs constructed by Jones [22], which were proved to have independence number $\alpha(G)=(n+1)/3$. Those graphs form a sequence $J_{1},J_{2},\ldots$ such that $J_{1}$ is the 5-cycle $(a_{1},b_{1},c_{1},d,e)$, and such that $J_{i+1}$ is obtained from $J_{i}$ by adding three vertices $a_{i+1},b_{i+1},c_{i+1}$ such that $N(a_{i+1})=\{b_{i},b_{i+1}\}$, $N(b_{i+1})=\{a_{i+1},c_{i+1}\}$, and $N(c_{i+1})=\{a_{i},c_{i},b_{i+1}\}$ (see Figure 8). It is already known that those graphs are planar, and it does only remain to show that they belong to CBU. Let us do so by exhibiting a homogeneous arc labeling $\ell$. This labeling is such that for any $i\geq 1$ we orient the edges $a_{i}b_{i}$ and $b_{i}c_{i}$ toward $b_{i}$, we orient the edges $x_{i}y_{i+1}$, for $x,y\in\{a,b,c\}$, from $x_{i}$ towards $y_{i+1}$, and we set $\ell(a_{i}b_{i})=\ell(a_{i}c_{i+1)}=2i$, $\ell(c_{i}b_{i})=\ell(c_{i}c_{i+1)}=2i$, and $\ell(b_{i}a_{i+1)}=2i+1$. By examining Figure 8 it is clear that this is a homogeneous arc labeling.
@@
A positive answer to this question, would give support to two conjectures. Let $\mathcal{P}_{g\geq 5}$ be the set of planar graph with girth at least five, and let $\mathcal{P}^{f}_{g\geq 4}$ be the set of planar graph with girth at least four, where every 4-cycle bounds a face. Clearly $\mathcal{P}_{g\geq 5}\subsetneq\mathcal{P}^{f}_{g\geq 4}$, since these classes avoid Jones graphs it is conjectured that graphs in $\mathcal{P}_{g\geq 5}$, or more generally graphs in $\mathcal{P}^{f}_{g\geq 4}$, have fractional chromatic number at most $c$, for some $c<3$ [13, 15]. However, our problem is not a sub-case of these conjectures (as $K_{2,t}$ belongs to $2$-CBU $\setminus\ \mathcal{P}^{f}_{g\geq 4}$), nor a super-case (as $\mathcal{P}_{g\geq 5}\setminus 2$-CBU is not empty, by Theorem 5.18).
@@ 9 Computational hardness for many problems
We have seen (c.f. Theorem 4.9, Corollary 4.11, and Corollary 5.12) that many 1-subdivided graphs belong to CBU, or even to 3- or 4-CBU. For $(\geq 2)$-subdivided graphs, the picture is even simpler.
Theorem 9.58.
@@
Proof 9.59.
Let us denote $v_{1},\ldots,v_{n}$ the vertices of $G$, and let $m=|E(G)|$. To construct a CBU representation for any $(\geq 2)$-subdivision, we start by assigning each vertex $v_{i}$ to the box $[3i,3i+1]\times[n-i,n-i+1]\times[0,2m]$. Then consider each edge $e$ of $G$ in any given order. For the $k^{\text{th}}$ edge $e$ assume it links $v_{i}$ and $v_{j}$, for some $i<j$, and assume $e$ is replaced by the path $(v_{i},u_{1},\ldots,u_{r},v_{j})$ for some $r\geq 2$. Here, $u_{1}$ is assigned to $[3i+1,3i+2]\times[n-j,n-i+1]\times[2k-1,2k]$, while the vertices $u_{\ell}$ with $2\leq\ell\leq r$ are assigned to $[3i+2+(\ell-2)(3j-3i-2)/(r-1),3i+2+(\ell-1)(3j-3i-2)/(r-1)]\times[n-j,n-j+1]\times[2k-1,2k]$ (see Figure 9). One can easily check that the obtained representation is a 3-CBU representation of the subdivided graph.
@@
Proof 9.61.
For Minimum Feedback Vertex Set and Cutwidth, this follows from the fact that these problems are NP-hard, and that for any instance, subdividing an edge does not change the solution. For Maximum Cut, it follows from its APX-hardness and the fact that the maximum cut of a graph $G$ and its 2-subdivision $G_{2\text{-sub}}$ verify $mc(G)=mc(G_{2\text{-sub}})-2|E(G)|$ and $3|E(G)|/2=|E(G_{2\text{-sub}})|/2\leq mc(G_{2\text{-sub}})\leq|E(G_{2\text{-sub}})|=3|E(G)|$. The other problems are shown APX-hard even when restricted to $6$-subdivided graphs [6].
@@
Proof 9.63.
As these problems belong to NP, it remains to show that they are NP-hard for 2-CBU graphs. Let us first show that the induced subgraphs of grids (so called grid graphs) belong to 2-CBU. Consider the $n\times n$ grid $G$ such that $V(G)=\{1,\ldots,n\}\times\{1,\ldots,n\}$, and such that the neighbors of any vertex $(i,j)$ are $\{(i,j-1)(i-1,j),(i,j+1),(i+1,j)\}\cap\{1,\ldots,n\}\times\{1,\ldots,n\}$. Since it suffices to delete some boxes to obtain an induced subgraph, the claim follows by constructing a 2-CBU representation for any such grid $G$. This construction is obtained by mapping any vertex $(i,j)$ to the box $[i+j-1,i+j]\times[2i-2j,2i-2j+3]$ (see Figure 10). As Domination [7], Hamiltonian Path, and Hamiltonian cycle [21] are NP-hard for grid graphs, those problems are NP-hard for 2-CBU graphs.
@@
For the problems Maximum Independent Set and Minimum Vertex Cover, we have to consider a variant of grid graphs, the graph $R^{\prime}(n_{1},n_{2})$ depicted in Figure 11, and it is easy to see how to modify the construction above in order to obtain a 2-CBU representation of this type of graphs. Again, this implies that every induced subgraph of such a graph belongs to 2-CBU. As the problems Maximum Independent Set and Minimum Vertex Cover are NP-hard for this class (see the proof of Theorem 10 in [19]), those problems are NP-hard for 2-CBU graphs
$(i-1,j)$$(i,j-1)$$(i+1,j)$$(i,j+1)$$(i-1,j+1)$$(i+1,j-1)$